OpenStudy (anonymous):

log_3(x+34)=log_3(x+16)=5

OpenStudy (bahrom7893):

Did u mean: log_3(x+34)+log_3(x+16)=5?

OpenStudy (cwrw238):

if so log_3 (x+34)(x+ 16) = 5 (x+34)(x+ 16) = 3^5 solve this and ur home